# Channel Knowledge Map-assisted Dual-domain Tracking and Predictive Beamforming for High-Mobility Wireless Networks

### v1

@@ VII-B Tracking Performance
Figure 5: Tracking performance for the proposed dual-domain framework versus the time index with $\sigma_{z}^{2}=10^{-9}\mathrm{\leavevmode\nobreak\ W}$ and $p_{\text{blk}}=0.15$.
Figure 6: Tracking performance for the proposed dual-domain framework versus the time index with $\sigma_{z}^{2}=10^{-6}\mathrm{\leavevmode\nobreak\ W}$ and $p_{\text{blk}}=0.15$.
(a) $\sigma_{z}^{2}=10^{-9}\mathrm{\leavevmode\nobreak\ W}$
(b) $\sigma_{z}^{2}=10^{-6}\mathrm{\leavevmode\nobreak\ W}$
@@
Figure 8: Multipath AoA tracking performance with $\sigma_{z}^{2}=10^{-6}\mathrm{\leavevmode\nobreak\ W}$.
@@
Fig. 5 illustrates the dual-domain tracking performance in terms of root mean squared error (RMSE) for the vehicle position and the absolute error for AoA estimation for path 1. In contrast to the baseline scheme, the proposed dual-domain EKF tracking framework ensures continuous target tracking and achieves a higher tracking accuracy even during the prolonged LoS-absent period. This is because our proposed scheme comprehensively models of the beam transition conditions and effectively utilizes the CKM to extract NLoS path information for positioning. Moreover, it can be observed that all the position tracking RMSE curves first increase and then decrease/stabilize with the evolution of time. This behavior is due to that the tracking accuracy of EKF algorithm requires several iterations to stabilize. Additionally, when path 1 is randomly blocked, the AoA of the strongest NLoS path changes unpredictably leading to beam misalignment and spikes in AoA tracking error indicated by asterisks in Fig. 5. Fortunately, these spikes do not significantly affect overall position tracking accuracy, as the measurement models in the C-Domain for LoS-present and LoS-absent cases are different. Note that the multipath AoA estimation accuracy has not been provided in Fig. 5 since the baseline only tracks the AoA of the LoS path. In the first few time slots, the AoA of LoS path in the following time slots of the baseline scheme is larger than that of the proposed framework because the MAP-based update principle proposed in Section V-C effectively incorporates a priori information when echo signal measurements are unreliable or have low SNR. In contrast, large errors in both target and beam tracking for the baseline scheme arise with time slot, even reaching more than $10$ meters of RMSE of position tracking and $50$ degrees of path 1 AoA tracking error in the end. This is because for LoS-absent cases, the echo signal measurement is treated as meaningless in the baseline scheme in [4] and relies on the state evolution model to predict the position of the vehicle and AoA of LoS path in the following time slots. Since the state evolution model adopted may suffer from model mismatch, with the increase of the slot, both the target and beam tracking error scales up.
Fig. 6 considers a higher noise power to evaluate the performance of the proposed scheme. When the noise power is higher, a larger antenna array can improve tracking accuracy due to a higher BF gain, but they also increase the likelihood of beam misalignment, as shown in Fig. 6. For a larger antenna array with $N_{t}=N_{r}=64$, both target and beam tracking performance becomes worse when the vehicle is close to the RSU, specifically between $80-100$ slots, even for the cases where path 1 has not been blocked. In fact, the AoA of path 1 varies rapidly in these time slots and B-Domain state prediction becomes inaccurate, and thus predictive BF with a narrow beam for a larger antenna array is more susceptible to missing the moving target.
Note that in both Fig. 5 and Fig. 6, baseline methods fail to maintain accuracy tracking in the considered complex and high-mobility wireless network due to random path blockage. The proposed CKM-assisted dual-domain tracking framework outperforms the baseline for two key reasons. First, based on the C-Domain tracking results, the CKM provides additional a priori information for B-Domain tracking, which improves AoA tracking accuracy, even under poor channel conditions. Second, conventional EKF-based target tracking schemes, such as those in prior works [7, 5, 4, 43], are highly sensitive to model mismatch errors due to discrepancy between the assumed measurement model and the actual propagation environment. In contrast, the proposed framework utilizes the CKM to serve as the measurement function in the LoS-absent case in addition to the traditional geometric-based measurement model in the LoS-present case.
Furthermore, Fig. 7a and Fig. 7b illustrate the cumulative distribution function (CDF) of AoA tracking error for path 1 and path 2 for different configurations of $N_{t}$ and $N_{r}$, corresponding to Figs. 5 and 6, respectively. Due to the lower noise power $\sigma_{z}^{2}$, Fig. 7a consistently exhibits better AoA estimation accuracy than Fig. 7b for all paths and antenna configurations. In contrast, Fig. 7b reveals that a higher measurement noise power significantly impacts the AoA tracking accuracy, particularly for path 2. This degradation is most pronounced during non-stationary and unpredictable beam state transitions, where the AoA estimation relies solely on the current observed signal and thus is more sensitive to noise-induced distortions. Consistent with the observations in Fig. 6, increasing the number of antennas is not always beneficial for AoA tracking, particularly for path 2 (NLoS path) and the case with a higher noise power. As the NLoS path has a higher random blockage probability, a smaller antenna array with a wider predictive beam has a higher probability of covering the bursty changed AoA of path 2 and thus may provide a high-quality signal measurement in the B-Domain. It is important to note that $p_{\text{blk}}=0.15$ indicates that $85\%$ of the AoA estimates for path 1 are obtained with the assistance of a priori information and $72.25\%$ of the AoA estimates for path 2 are supported with a priori information. In both Figs. 7a and 7b, the AoA tracking error is lower than $10$ degrees with the assistance of a priori information. This indicates that incorporating a priori information significantly improves AoA tracking accuracy, even under low SNR conditions, especially for the NLoS path. This finding underscores the robustness of our proposed approach in maintaining accurate angle estimates, even in the presence of significant noise and high-speed moving target. Fig. 10 further corroborates the analysis above, demonstrating that while larger antenna arrays can provide a better AoA tracking accuracy in stationary or predictable cases, they are more sensitive to unexpected beam state transition in the propagation environment.
Fig. 10 further demonstrates that nonstationary and unpredictable transition $(III)$ degrades the AoA tracking performance compared to the other two types of beam state transition, particularly for NLoS path and the case with a larger antenna array. In fact, in the third type of beam state transition, without a prior information of channel and relying solely on the received signal, it is challenging to estimate the AoA for the NLoS path due to its much higher propagation loss than that of LoS path. Moreover, in this case, the RSU struggles to predict the beam and predictive BF may suffer from the beam misalignment with a higher probability for a larger antenna array.
VII-C Predictive BF Performance

### v2

@@ VII-B Tracking Performance
Figure 6: Tracking performance for the proposed dual-domain framework versus the time index with $\sigma_{z}^{2}=10^{-9}\mathrm{\penalty 10000\ W}$ and $p_{\text{blk}}=0.15$.
Figure 7: Tracking performance for the proposed dual-domain framework versus the time index with $\sigma_{z}^{2}=10^{-6}\mathrm{\penalty 10000\ W}$ and $p_{\text{blk}}=0.15$.
(a) $\sigma_{z}^{2}=10^{-9}\mathrm{\penalty 10000\ W}$
(b) $\sigma_{z}^{2}=10^{-6}\mathrm{\penalty 10000\ W}$
@@
Fig. 6 illustrates the dual-domain tracking performance in terms of root mean squared error (RMSE) for the vehicle position and the absolute error for AoA estimation for path 1. In contrast to the baseline scheme, the proposed dual-domain EKF tracking framework ensures continuous target tracking and achieves a higher tracking accuracy even during the prolonged LoS-absent period. This is because our proposed scheme comprehensively models of the beam transition conditions and effectively utilizes the CKM to extract NLoS path information for positioning. Moreover, it can be observed that all the position tracking RMSE curves first increase and then decrease/stabilize with the evolution of time. This behavior is due to that the tracking accuracy of EKF algorithm requires several iterations to stabilize. Additionally, when path 1 is randomly blocked, the AoA of the strongest NLoS path changes unpredictably leading to beam misalignment and spikes in AoA tracking error indicated by asterisks in Fig. 6. The AoA tracking errors are mostly concentrated near 0°, indicating that the B-Domain tracking maintains a high AoA estimation accuracy even under random blockages. In contrast, a large tracking error happens when the LoS path disappears due to random blockage and the induced path lies far from the predicted beam direction, even occasionally approaching 180°. This is because a sudden change in the B-Domain state leads to severe beam misalignment and a significantly reduced received SNR for the induced path. Fortunately, these spikes do not significantly affect overall position tracking accuracy, as the measurement models in the C-Domain for LoS-present and LoS-absent cases are different. Note that the multipath AoA estimation accuracy has not been provided in Fig. 6 since the baseline only tracks the AoA of the LoS path. In the first few time slots, the AoA of LoS path in the following time slots of the baseline scheme is larger than that of the proposed framework because the MAP-based update principle proposed in Section V-C effectively incorporates a priori information when echo signal measurements are unreliable or have low SNR. In contrast, large errors in both target and beam tracking for the baseline scheme arise with time slot, even reaching more than $10$ meters of RMSE of position tracking and $50$ degrees of path 1 AoA tracking error in the end. This is because for LoS-absent cases, the echo signal measurement is treated as meaningless in the baseline scheme in [53] and relies on the state evolution model to predict the position of the vehicle and AoA of LoS path in the following time slots. Since the state evolution model adopted may suffer from model mismatch, with the increase of the slot, both the target and beam tracking error scale up.
Fig. 7 considers a higher noise power to evaluate the performance of the proposed scheme. When the noise power is higher, a larger antenna array can improve tracking accuracy due to a higher BF gain, but they also increase the likelihood of beam misalignment, as shown in Fig. 7. For a larger antenna array with $N_{t}=N_{r}=64$, both target and beam tracking performance become worse when the vehicle is close to the RSU, specifically between 8010080-100 slots, even for the cases where path 1 has not been blocked. In fact, the AoA of path 1 varies rapidly in these time slots and B-Domain state prediction becomes inaccurate, and thus predictive BF with a narrow beam for a larger antenna array is more susceptible to missing the moving target.
Note that in both Fig. 6 and Fig. 7, baseline methods fail to maintain accurate tracking in the considered complex and high-mobility wireless network due to random path blockage. The proposed CKM-assisted dual-domain tracking framework outperforms the baseline for two key reasons. First, based on the C-Domain tracking results, the CKM provides additional a priori information for B-Domain tracking, which improves AoA tracking accuracy, even under poor channel conditions. Second, conventional EKF-based target tracking schemes, such as those in prior works [7, 21, 53, 43], are highly sensitive to model mismatch errors due to discrepancy between the assumed measurement model and the actual propagation environment. In contrast, the proposed framework utilizes the CKM to serve as the measurement function in the LoS-absent case in addition to the traditional geometric-based measurement model in the LoS-present case.
@@
Figure 9: Multipath AoA tracking performance with $\sigma_{z}^{2}=10^{-6}\mathrm{\penalty 10000\ W}$.
Furthermore, Fig. 8a and Fig. 8b illustrate the CDF of AoA tracking error for path 1 and path 2 for different configurations of $N_{t}$ and $N_{r}$, corresponding to Figs. 6 and 7, respectively. Due to the lower noise power $\sigma_{z}^{2}$, Fig. 8a consistently exhibits better AoA estimation accuracy than Fig. 8b for all paths and antenna configurations. In contrast, Fig. 8b reveals that a higher measurement noise power significantly impacts the AoA tracking accuracy, particularly for path 2. This degradation is most pronounced during non-stationary and unpredictable beam state transitions, where the AoA estimation relies solely on the current observed signal and thus is more sensitive to noise-induced distortions. Consistent with the observations in Fig. 7, increasing the number of antennas is not always beneficial for AoA tracking, particularly for path 2 (NLoS path) and the case with a higher noise power. As the NLoS path has a higher random blockage probability, a smaller antenna array with a wider predictive beam has a higher probability of covering the bursty changed AoA of path 2 and thus may provide a high-quality signal measurement in the B-Domain. It is important to note that $p_{\text{blk}}=0.15$ indicates that $85\%$ of the AoA estimates for path 1 are obtained with the assistance of a priori information and $72.25\%$ of the AoA estimates for path 2 are supported with a priori information. In both Figs. 8a and 8b, the AoA tracking error is lower than $10$ degrees with the assistance of a priori information. This indicates that incorporating a priori information significantly improves AoA tracking accuracy, even under low SNR conditions, especially for the NLoS path. This finding underscores the robustness of our proposed approach in maintaining accurate angle estimates, even in the presence of significant noise and high-speed moving target. Fig. 9 further corroborates the analysis above, demonstrating that while larger antenna arrays can provide a better AoA tracking accuracy in stationary or predictable cases, they are more sensitive to unexpected beam state transition in the propagation environment.
@@
Fig. 10 further demonstrates that non-stationary and unpredictable transition $(III)$ degrades the AoA tracking performance compared to the other two types of beam state transition, particularly for NLoS path and the case with a larger antenna array. In fact, in the third type of beam state transition, without a prior information of channel and relying solely on the received signal, it is challenging to estimate the AoA for the NLoS path due to its much higher propagation loss than that of LoS path. Moreover, in this case, the RSU struggles to predict the beam and predictive BF may suffer from the beam misalignment with a higher probability for a larger antenna array.
VII-C Predictive BF Performance